# A spring–block theory of feature learning in deep neural networks

### v2

@@ VI.3 Second order Langevin equation
$f(v)=\begin{cases}\mu_{\rightarrow}&\text{if~{}}v>0\\ -\mu_{\leftarrow}&\text{if~{}}v<0.\end{cases}$ (13)
@@
In Eq. (12), $W_{\ell}(t)$ represents the Wiener process with $\varepsilon$ controlling the amount of the noise (temperature parameters). To ensure convergence in this formulation we have to decay the noise; we set $\varepsilon(t)=\varepsilon_{0}e^{-\tau t}$. The friction at zero speed $f(0)$ is defined similarly to the static friction $f_{\ell}$ in Eq. (10), but with $F_{\ell}=k(\bm{{L}}x)_{\ell}$ without considering noise. This dynamics can be solved by standard SDE integration. It results in a similar phase diagram at convergence, as shown in Fig. S14, although the fluctuations do not appear as similar as with the formulation in the main text.
@@
Figure S14: Phase diagram of spring block system with Langevin equation (12) for friction $\mu_{\rightarrow}$ vs. noise level $\epsilon_{0}$. We set $k=0.1,\gamma=1,\tau=0.002,\mu_{\leftarrow}=0.016$ and $L=7$. Red load curves ($D_{\ell}=y-x_{\ell}$) are recorded at $t=300$; the shading corresponds to the elastic potential energy.

### v4

@@ F.3 Second order Langevin equation
$f(v)=\begin{cases}\mu_{\rightarrow}&\text{if~{}}v>0\\ -\mu_{\leftarrow}&\text{if~{}}v<0.\end{cases}$ (S5)
@@
In Eq. (S4), $W_{\ell}(t)$ represents the Wiener process with $\varepsilon$ controlling the amount of the noise (temperature parameters). To ensure convergence in this formulation we have to decay the noise; we set $\varepsilon(t)=\varepsilon_{0}e^{-\tau t}$. The friction at zero speed $f(0)$ is defined similarly to the static friction $f_{\ell}$ in Eq. (S2), but with $F_{\ell}=k(\bm{{L}}x)_{\ell}$ without considering noise. This dynamics can be solved by standard SDE integration. It results in a similar phase diagram at convergence, as shown in FIG. S18, although the fluctuations do not appear as similar as with the formulation in the main text.
@@
Figure S18: Phase diagram of spring block system with Langevin equation (S4) for friction $\mu_{\rightarrow}$ vs. noise level $\epsilon_{0}$. We set $k=0.1,\gamma=1,\tau=0.002,\mu_{\leftarrow}=0.016$ and $L=7$. Red load curves ($D_{\ell}=y-x_{\ell}$) are recorded at $t=300$; the shading corresponds to the elastic potential energy.
Appendix G S7. Scaling and quantity of noise sources
In FIG. 1, we show that four “noise” sources—data noise, learning rate, dropout, and batch size—exhibit similar phase diagrams. Although it is difficult to precisely quantify these different factors in deep neural networks in an integrated manner, their scaling (whether logarithmic, linear, or otherwise on the x-axis of the phase diagram) can still be verified using known results on stochastic modified equations [69]. Specifically, since SGD can be approximated (in a certain weak sense) by a stochastic differential equation (SDE), by analyzing the structure of the drift and diffusion terms in this SDE, we can understand the role and scaling of different types of noise. Here, we use the learning rate as a reference to examine how other hyperparameters should be adjusted when varying the learning rate to keep the SDE unchanged. For example, standard SGD can be written as
$\bm{\theta}_{k+1}=\bm{\theta}_{k}-\eta\nabla_{\bm{\theta}}\left(\frac{1}{|\mathcal{B}_{k}|}\sum_{i\in\mathcal{B}_{k}}l\left(\bm{\theta}_{k};(\bm{{x}}_{i},\bm{{y}}_{i})\right)\right),$
where $\eta$ is the learning rate, and the mini-batch $\mathcal{B}_{k}$ is sampled uniformly from a large dataset. For a typical loss function, neural network, and data distribution $(\bm{{x}}_{i},\bm{{y}}_{i})$, we can treat $\nabla{\bm{\theta}}l(\bm{\theta};(\bm{{x}}_{i},\bm{{y}}_{i}))$ as a random variable with mean $\bm{{g}}(\bm{\theta})$ and covariance $\bm{{C}}(\bm{\theta})$. For a large batch size $B=|\mathcal{B}_{k}|$, by the central limit theorem the random fluctuations are approximately normal $\bm{{\xi}}^{C}_{k}\sim\mathcal{N}(0,\bm{{C}}(\bm{\theta}_{k}))$, so that the SGD steps (approximately) read
$\bm{\theta}_{k+1}=\bm{\theta}_{k}-\eta\bm{{g}}(\bm{\theta}_{k})+\frac{\eta}{\sqrt{B}}\bm{{\xi}}^{C}_{k}.$
This can be recognized as the Euler–Maruyama discretization of the following SDE
$\mathrm{d}\bm{\theta}=-\bm{{g}}(\bm{\theta})\mathrm{d}t+\sqrt{\frac{\eta}{B}}\bm{{C}}^{\frac{1}{2}}(\bm{\theta})\mathrm{d}\bm{{W}}(t),$
where $\bm{{W}}(t)$ is a standard Brownian motion. In this case, if we plot the learning rate $\eta$ on a logarithmic scale, the batch size $B$ should also be plotted on a logarithmic scale. This relationship has been observed and studied in earlier works, e.g. [70, 71, 72].
Dropout in deep neural networks (DNNs) is more complex, as the stochasticity occurs at each layer and accumulates across the depth. A recent work [73] analyzed this problem in a two-layer network with MSE loss. The authors approximate the trajectories of full batch GD with dropout by the following SDE:
$\mathrm{d}\bm{\theta}$ $=-\left(\bm{{g}}_{1}(\bm{\theta})+\frac{q}{1-q}\bm{{g}}_{2}(\bm{\theta})\right)\mathrm{d}t$
$+\sqrt{\eta}\left(\frac{q}{1-q}\bm{{C}}_{1}(\bm{\theta})+\frac{q}{(1-q)^{2}}\bm{{C}}_{2}(\bm{\theta})\right)^{1/2}\mathrm{d}\bm{{W}}(t),$
where $q$ is the probability of setting activations to zero. For small dropout, as $q\to 0$, we can approximate it as $\mathrm{d}\bm{\theta}\approx-\bm{{g}}_{1}(\bm{\theta})\mathrm{d}t+\sqrt{\eta q}(\bm{{C}}_{1}+\bm{{C}}_{2})^{\frac{1}{2}}\mathrm{d}\bm{{W}}(t).$ It implies that if we plot the learning rate $\eta$ on a logarithmic scale, the dropout ratio $q$ should also be plotted on a logarithmic scale.
In the data noise experiments, we randomly reset a fraction $p$ of the labels while simultaneously adding Gaussian noise with variance $p^{2}$ to the input data. When $p$ is away from zero, the noise introduced by label corruption is significantly larger than the noise from the input data. Therefore, we focus primarily on analyzing label noise. Since cross-entropy loss is biased and scale-sensitive [74], most analytical studies on label noise are based on regression settings with unbiased Gaussian noise, e.g., [75, 76]. However, we can still analyze a simple setting to estimate the scale based on the method we used previously.
To simplify the problem, we study a single data point $\bm{{x}}$ with label $1$ and a one-layer network $f_{m}(x)=\langle\bm{{w}}_{m},\bm{{x}}\rangle$ with $m=1,\dots,M$, corresponding to $M$ classes. The cross-entropy loss with softmax can be computed as
=$\mathcal{L}=\begin{cases}-\log\frac{\exp\langle\bm{{w}}_{1},\bm{{x}}\rangle}{\sum_{m}\exp\langle\bm{{w}}_{m},\bm{{x}}\rangle}&\quad\text{w. p. $1-p+\frac{p}{M}$,}\\ -\log\frac{\exp\langle\bm{{w}}_{j},\bm{{x}}\rangle}{\sum_{m}\exp\langle\bm{{w}}_{m},\bm{{x}}\rangle}&\quad\text{for $j>1$, w. p. $\frac{p}{M}$.}\end{cases}$
We can then compute the mean and variance of each gradient descent step and approximate the GD by the stochastic modified equations SDE as mentioned in [69, 73]. After rescaling the loss function, and assuming the number of classes is large, $M\gg 1$, we obtain a simple formula:
$\mathrm{d}\bm{{w}}_{1}=\left(1-\frac{1}{1-p}t_{1}(\bm{{w}})\right)\bm{{x}}+\sqrt{\eta\frac{p}{1-p}}\left(\bm{{x}}\bm{{x}}^{T}\right)^{\frac{1}{2}}\mathrm{d}\bm{{W}}(t),$
where $t_{1}=\frac{\exp\langle\bm{{w}}_{1},\bm{{x}}\rangle}{\sum_{m}\exp\langle\bm{{w}}_{m},\bm{{x}}\rangle}$. The stochasticity of the remaining weights $\bm{{w}}_{m}$ for $m>1$ is negligible for large $M$. We also note that $t_{1}$ is small at initialization, so to obtain a consistent phase diagram, it is better to set $\frac{p}{1-p}$ in logarithmic scale when using $\eta$ in logarithmic scale. Noting that we are not only interested in small $p$, we can leverage the well-known approximation of the logit function:
$\mathrm{logit}(x)=\log\left(\frac{x}{1-x}\right)\approx 4(x-0.5)\quad\text{for}x\text{ near }0.5,$
which leads to a linear scale of x-axis in the first panel in FIG. 1.
Figure S19: Phase diagrams of DNN training load curves (solid red) and test load curves (dashed red) of the experiments in FIG. 1 shown together.
Figure S20: Phase diagram of test load curve of the experiments in FIG. 1.